# Bessel potentials and Green functions on pseudo-Euclidean spaces

### v1

@@ Abstract
We review properties of Bessel potentials, that is, inverse Fourier transforms of (regularizations of) $\frac{1}{(m^{2}+p^{2})^{\frac{\mu}{2}}}$ on a pseudoEuclidean space with signature $(q,d-q)$. We are mostly interested in the Lorentzian signature $(1,d-1)$, and the case $\mu=2$, related to the Klein-Gordon equation $(-\Box+m^{2})f=0$. We analyze properties of various “two-point functions”, which play an important role in Quantum Field Theory, such as the retarded/advanced propagators or Feynman/antiFeynman propagators. We consistently use hypergeometric functions instead of Bessel functions, which makes most formulas much more transparent. We pay attention to distributional properties of various Bessel potentials. We include in our analysis the “tachyonic case”, corresponding to the “wrong” sign in the Klein-Gordon equation.
@@ 1 Introduction
on the Euclidean space ${\mathbb{R}}^{d}$.
@@
so the case $m>0$ reduces to $m=1$. $G_{\mu,1}(x)$ can be expressed in terms of the Macdonald function, one of solutions of the modified Bessel equation. Therefore, $G_{\mu,1}(x)$ is often called the Bessel potential of order $\mu$.
The function $G_{\mu,0}(x)$ is called the Riesz potential of order $\mu$.
The case $\mu=2$ is probably the most important. In this case we will usually omit $\mu$ from the notation, setting $G_{m}(x):=G_{2,m}(x)$ and obtaining the Green function of the inhomogeneous Helmholtz equation
@@
The definition (1.1) is usually no longer correct, since $\frac{1}{(m^{2}+p^{2})^{\frac{\mu}{2}}}$ may fail to be locally integrable, and hence may not define a tempered distribution. A possible generalization of (1.1) is then the pair of functions
@@
The Lorentzian case is especially interesting and rich. This is related to the fact that the Minkowski space ${\mathbb{R}}^{1,d-1}$ can be equipped with a causal structure and the set $p^{2}+m^{2}=0$ has two connected components. Therefore, the distributions
@@
are invariant wrt the orthochronous Lorentz transformations. Remarkably, $G_{\mu,m}^{\vee/\wedge}$ is supported in the forward, resp. backward cone. Therefore, $G_{\mu,m}^{\vee}$ is called the forward (or retarded), and $G_{\mu,m}^{\wedge}$ the backward (or advanced) Bessel potential.
@@
The Klein-Gordon equation possesses many Green functions. Among them, we have the Feynman and antiFeynman Green functions given by the formulas (1.6) and (1.7) with $\mu=2$. Another distinguished pair consists of the retarded (or forward) Green function $G^{\vee}$ and the advanced (or backward) Green function $G^{\wedge}$. They are defined by demanding that the support of $G^{\vee}$, resp. $G^{\wedge}$ is contained in the forward, resp. backward cone. They are given by (1.8) and (1.9) with $\mu=2$.
The Feynman, anti-Feynman, forward, and backward Green functions of the Klein-Gordon equation have important applications in physics, especially in classical and quantum field theory. The forward and backward Green functions can be used to express the Cauchy problem. The Feynman, resp. anti-Feynman Green functions express the time-ordered, resp. anti-time-ordered vacuum expectation values of fields in quantum field theory. Note the identity
@@
Remarkably, all four basic Green functions, Feynman $G_{m}^{\mathrm{F}}$, anti-Feynman $G_{m}^{\overline{\mathrm{F}}}$, forward $G_{m}^{\vee}$, and backward $G_{m}^{\wedge}$, can be defined in the tachyonic case. For the Feynman and anti-Feynman Green functions we can still use the formulas (1.6) and (1.7), where $m^{2}$ is replaced with $-m^{2}$. Their interpretation in terms of the vacuum expectation values is however lost since the tachyonic theory has no vacuum. For the forward and backward Green functions, we cannot use the formulas (1.8) and (1.9). In fact, the set $p^{2}-m^{2}=0$ is now connected, and cutting it with ${\rm sgn}p^{0}$ is no longer invariant. Nevertheless, one can use the analytic continuation in $m$ to uniquely define Green functions with correct support properties also in the tachyonic case. Interestingly, the identity (1.13) is no longer true in the tachyonic case.
@@
The difference of two Green functions is a solution of the homogeneous Helmholtz/Klein-Gordon equation. Certain distinguished solutions are important for physics applications. In the Lorentzian case, we have the Pauli-Jordan propagator, for $m^{2}\geq 0$, and the positive frequency and the negative frequency two-point functions. We illustrate applications of distinguished solutions to the Helmholtz/Klein-Gordon equation by computing averages of plane waves over the sphere (in the Euclidean case), as well as over the hyperbolic and de Sitter space (in the Lorentzian case).
@@
Let us say a few words about the history of Bessel potentials. The name Bessel potentials was introduced in the 60s by Aronszajn and Smith, who studied them in the Euclidean case in [1]. Around the same time, they were also investigated by Calderon [2]. Bessel potentials are frequently viewed in the literature as smoothed versions of Riesz potentials (see, for example, [3] where they are defined using the integral formula (2.5)). For a comprehensive treatment of (Euclidean) Bessel potentials, we refer the reader to [1], where many properties of Bessel potentials are exhaustively studied.
@@
The Lorentzian versions of Bessel potentials, typically in dimension 1+3, often appear in the literature on Quantum Field Theory. For instance, the famous textbooks by Björken-Drell [4] and by Bogoliubov– Shirkov [5] contain appendices devoted to distinguished Green functions and solutions of the Klein-Gordon equation in the physical dimension 1+3. They carry various names. For instance, often the term Green function is replaced by propagator, etc.
@@
Formulas for Bessel potentials in various signatures are available in collections of integrals such as [6] and [7]. In chapter III.2 of [8] one can find Fourier transforms of powers of quadratic forms with any signature, including the formula (4.5) of the general case studied in this paper. Although there exists a large literature about Bessel potentials, our presentation contains several new points, which we have not seen in the literature and believe are important.
@@
which can be called the ${}_{0}F_{1}$ hypergeometric equation. Equation (1.15) has two singular points: 00 and $\infty$. The singularity at 00 is regular (Fuchsian), and the solution obtained by the well-known Frobenius is the ${}_{0}F_{1}$ hypergeometric function, which we denote $F_{\alpha}$. We usually prefer its Olver normalized version ${\bf F}_{\alpha}:=\frac{F_{\alpha}}{\Gamma(\alpha+1)}$, closely related to the Bessel function, both standard and modified.
Another standard solution of the ${}_{0}F_{1}$ equation, corresponding to the irregular singularity at $\infty$, is the function that we denote $U_{\alpha}$, This function is perhaps less known. Up to a coefficient, it coincides with the Meijer $G$-function $G_{0,2}^{2,0}(-;0,-\alpha;z)$. Function $U_{\alpha}$ is closely related to the Macdonald and Hankel functions.
In our opinion, it is much more convenient and transparent to use ${\bf F}_{\alpha}$ and $U_{\alpha}$ to express Bessel potentials, rather than to use functions from the Bessel family, as it is done in the conventional treatment of this topic. The corresponding formulas are simpler and more transparent. This is especially visible when we consider non-Euclidean signatures, where the formulas involve analytic continuation across two branches and an irregular distribution at the junction of these branches. The ${\bf F}_{\alpha}$ and $U_{\alpha}$ functions are also convenient to see the transition from the Minkowski space to the deSitter and the universal cover of the AntideSitter space, as discussed in [9]. Therefore, in our paper, we treat these functions as basic elements of our description of Bessel potentials.
@@
In our discussion we cover not only the massive and massless case, but also the tachyonic case. This case is quite curious, even though usually ignored in the physics literature.
Finally, in our treatment, we pay special attention to the distributional character of Bessel potentials. This is unproblematic in the Euclidean signature, where Bessel potentials are given by (locally) integrable functions. This is not the case in non-Euclidean signatures. In particular, it is interesting to look at the functions ${\bf F}_{\alpha}$ and $U_{\alpha}$ as defining distributions on the real line. With this interpretation in mind, well-known identities have to be reformulated see e.g. (2.47).
@@ 2.1 The ${}_{0}F_{1}$ equation
In this section, we mostly follow [10], see also [11].
@@
$(z{\rm\partial}_{z}^{2}+c{\rm\partial}_{z}-1)v(z).$ (2.1)
If $c\neq 0,-1,-2,\dots$, then the only solution of the ${}_{0}F_{1}$ equation $=1$ at 0 is called the ${}_{0}F_{1}$ hypergeometric function:
$F(c;z):=\sum_{j=0}^{\infty}\frac{1}{(c)_{j}}\frac{z^{j}}{j!}.$
It is defined for $c\neq 0,-1,-2,\dots$. Sometimes it is more convenient to consider the function
@@
As $|z|\to\infty$ and $|\arg z|<\pi/2-\epsilon$, we have
@@
$U_{\alpha}$ is a unique solution of (1.15) with this property.
@@ 2.3 Relationship to Meijer G-functions
Solutions of hypergeometric equations ${}_{p}F_{q}$ can be expressed in terms of Meijer $G$-functions [12]. In particular, the ${}_{0}F_{1}$ equation can be solved by
@@
Here, the contour $L_{1}$ goes from +$\infty$+ ∞ to +$\infty$+ ∞ and encircles ${\mathbb{N}}_{0}$, and the contour $L_{2}$ also goes from +$\infty$+ ∞ to +$\infty$+ ∞ and encircles ${\mathbb{N}}_{0}\cup({\mathbb{N}}_{0}-\alpha)$. Computing the residues and using the connection formula (2.8) we obtain
@@ 2.4 Relationship to Bessel functions
In the literature, the ${}_{0}F_{1}$ equation is seldom used. Much more frequent is the modified Bessel equation, which is equivalent to the ${}_{0}F_{1}$ equation:
@@
Even more frequent is the (standard) Bessel equation:
@@
The ${}_{0}F_{1}$ function is also seldom used. Instead, one uses the modified Bessel function and, even more frequently, the Bessel function:
@@ 2.5 ${\bf F}_{\alpha}$ and $U_{\alpha}$ functions as distributions
where the right-hand side should be understood as the limit in the distributional sense. Note that for ${w}\neq 0$ these distributions are regular and given by analytic functions:
@@
(Note that both sides of (2.47) are zero for $w>0$). It is easy to see that for $\alpha\not\in{\mathbb{N}}$ (2.47) is a correct distributional distributional identity, where the lhs is the product of the distribution $w_{-}^{-\alpha}$ and of the smooth function ${\bf F}_{-\alpha}(w)$, whereas the rhs is a linear combination of distributions defined in (2.42). (2.47) can be decomposed into a singular and regular part as follows:

### v2

@@ Abstract
We review properties of Bessel potentials, that is, inverse Fourier transforms of (regularizations of) $(m^{2}+p^{2})^{-\frac{\mu}{2}}$ on a pseudoEuclidean space with signature $(q,d-q)$. We are mostly interested in the Lorentzian signature $(1,d-1)$, and the case $\mu=2$, related to the Klein-Gordon equation $(-\Box+m^{2})f=0$. We analyze properties of various “propagators”, which play an important role in Quantum Field Theory, such as the retarded/advanced propagators or Feynman/antiFeynman propagators. We consistently use hypergeometric functions instead of Bessel functions, which makes most formulas much more transparent. We pay attention to distributional properties of various Bessel potentials. We include in our analysis the “tachyonic case”, corresponding to the “wrong” sign in the Klein-Gordon equation.
Keywords: Bessel potential, Riesz potential, Klein-Gordon equation, Minkowski space.
@@ 1 Introduction
on the Euclidean space ${\mathbb{R}}^{d}$. Note that that $G_{\mu,m}(x-y)$ can be interpreted as the integral kernel of the operator $(m^{2}-\Delta)^{-\frac{\mu}{2}}$.
@@
so the case $m>0$ reduces to $m=1$. $G_{\mu,1}(x)$ can be expressed in terms of the Macdonald function, one of solutions of the modified Bessel equation. Therefore, $G_{\mu,1}(x)$ is often called the Bessel potential of order $\mu$. The function $G_{\mu,0}(x)$ is called the Riesz potential of order $\mu$.
It is remarkable that the theory of Bessel potentials is very similar for all $\mu>0$. However, the case $\mu=2$ is probably the most important. In this case we will usually omit $\mu$ from the notation, setting $G_{m}(x):=G_{2,m}(x)$, and obtaining the Green function of the inhomogeneous Helmholtz equation
@@
The definition (1.1) is usually no longer correct for $m^{2}\in{\mathbb{R}}$, since $\frac{1}{(m^{2}+p^{2})^{\frac{\mu}{2}}}$ may fail to be locally integrable, and hence may not define a tempered distribution. It still works for complex non-real $m^{2}$. A possible pair of generalizations of (1.1) to $m^{2}$ real is the pair of functions, which correspond to the limits from above and below:
@@
(1.6) and (1.7) have an obvious interpretation as boundary values of integral kernels of appropriate functions of the pseudoLaplacian
$\Box:=-\partial_{1}^{2}\cdots-\partial_{q}^{2}+\partial_{q+1}^{2}+\partial_{d}^{2}.$ (1.8)
@@
The Lorentzian case is especially interesting and rich. This is related to the fact that the Minkowski space ${\mathbb{R}}^{1,d-1}$ can be equipped with a causal structure and the set $p^{2}+m^{2}=0$ has two connected components. Therefore, besides $G_{\mu,m}^{\mathrm{F}/\overline{\mathrm{F}}}$, we can introduce the distributions
@@
which are invariant wrt orthochronous Lorentz transformations. Remarkably, $G_{\mu,m}^{\vee/\wedge}$ is supported in the forward, resp. backward cone. Therefore, $G_{\mu,m}^{\vee}$ is called the forward (or retarded), and $G_{\mu,m}^{\wedge}$ the backward (or advanced) Bessel potential.
@@
The Klein-Gordon equation possesses many Green functions. Among them, we have the Feynman and antiFeynman Green functions given by the formulas (1.6) and (1.7) with $\mu=2$. Another distinguished pair consists of the retarded (or forward) Green function and the advanced (or backward) Green function, defined by demanding that their support is contained in the forward, resp. backward cone. For $m^{2}\geq 0$ the retarded Green function is given by (1.9) and the advanced Green function by (1.10) with $\mu=2$.
The Feynman, anti-Feynman, forward, and backward Green functions of the Klein-Gordon equation have important applications in physics, especially in classical and quantum field theory. The forward and backward Green functions can be used to express the Cauchy problem. The Feynman, resp. anti-Feynman Green functions express the time-ordered, resp. anti-time-ordered vacuum expectation values of fields in quantum field theory. Importantly, they satisfy the identity
@@
Remarkably, all four basic Green functions, Feynman $G_{m}^{\mathrm{F}}$, anti-Feynman $G_{m}^{\overline{\mathrm{F}}}$, forward $G_{m}^{\vee}$, and backward $G_{m}^{\wedge}$, can be defined in the tachyonic case. For the Feynman and anti-Feynman Green functions we can still use the formulas (1.6) and (1.7), where $m^{2}$ is replaced with $-m^{2}$. Their interpretation in terms of the vacuum expectation values is however lost, since the tachyonic theory has no vacuum state. (In particular, in the tachyonic case we do not have a counterpart of the positive/negative frequency Green functions (5.40)). The forward and backward Green functions are defined by their support properties. For them we cannot use the formulas (1.9) and (1.10). In fact, the set $p^{2}-m^{2}=0$ is now connected, and cutting it with ${\rm sgn}p^{0}$ is no longer invariant. Nevertheless, one can use the analytic continuation in $m$ to uniquely define Green functions with correct support properties also in the tachyonic case. We point out that the identity (1.14) is no longer true in the tachyonic case.
@@
The difference of two Green functions is a solution of the homogeneous Helmholtz/Klein-Gordon equation. Certain distinguished solutions are important for physics applications. In the Lorentzian case, we have the Pauli-Jordan propagator; for $m^{2}\geq 0$ also the positive frequency and the negative frequency two-point functions. We illustrate applications of distinguished solutions to the Helmholtz/Klein-Gordon equation by computing averages of plane waves over the sphere (in the Euclidean case), as well as over the hyperbolic and de Sitter space (in the Lorentzian case).
@@
Let us say a few words about the history of Bessel potentials. The name Bessel potentials was introduced in the 60s by Aronszajn and Smith, who studied them in the Euclidean case in [1]. Around the same time, they were also investigated by Calderon [2]. Bessel potentials are frequently viewed in the literature as smoothed versions of Riesz potentials (see, for example, [3] where they are defined using the integral formula (2.5)). They are often used to define Bessel potential spaces that generalize standard Sobolev spaces (see [4]), and the idea to use Bessel kernels is due to Deny [5]. For a comprehensive treatment of (Euclidean) Bessel potentials, we refer the reader to [1], where many properties of Bessel potentials are exhaustively studied.
@@
The Lorentzian versions of Bessel potentials, typically in dimension 1+3, often appear in the literature on Quantum Field Theory. They are ingredients of formulas for scattering amplitudes based on Feynman diagrams and on the Epstein-Glaser approach [6, 7]. The famous textbooks by Björken-Drell [8] and by Bogoliubov– Shirkov [9] contain appendices devoted to distinguished Green functions and solutions of the Klein-Gordon equation in the physical dimension 1+3. They carry various names. For instance, often the term Green function is replaced by propagator, etc.
@@
Formulas for Bessel potentials in various signatures are known and are available in collections of integrals such as [10] and [11]. In chapter III.2 of [12] one can find Fourier transforms of powers of quadratic forms with any signature, including the formula (4.5) of the general case studied in this paper. Although there exists a large literature about Bessel potentials, our presentation contains several new points, which we have not seen in the literature and believe are important.
@@
which can be called the ${}_{0}F_{1}$ hypergeometric equation. Equation (1.16) has two singular points: 00 and $\infty$. The singularity at 00 is regular (Fuchsian), and the solution obtained by the well-known Frobenius method is the ${}_{0}F_{1}$ hypergeometric function, which we denote $F_{\alpha}$. We usually prefer its Olver normalized version ${\bf F}_{\alpha}:=\frac{F_{\alpha}}{\Gamma(\alpha+1)}$, closely related to the Bessel function, both standard and modified.
Another standard solution of the ${}_{0}F_{1}$ equation, corresponding to the irregular singularity at $\infty$, is the function that we denote $U_{\alpha}$, This function is perhaps less known. Up to a coefficient, it coincides with the Meijer $G$-function $G_{0,2}^{2,0}(-;0,-\alpha;z)$. The function $U_{\alpha}$ is closely related to the Macdonald and Hankel functions.
In our paper, we treat ${\bf F}_{\alpha}$ and $U_{\alpha}$ functions as basic elements of our description of Bessel potentials. In our opinion, they are much more convenient for this purpose, rather than functions from the Bessel family, as it is done in the conventional treatment of this topic. The corresponding formulas are simpler and more transparent. This is especially visible when we consider non-Euclidean signatures, where the formulas involve analytic continuation across two branches and an irregular distribution at the junction of these branches. The ${\bf F}_{\alpha}$ and $U_{\alpha}$ functions are also convenient to see the transition from the Minkowski space to the deSitter and the universal cover of the AntideSitter space, as discussed in [13]. In fact, on the Minkowski space retarded/advanced and Feynman/anti-Feynman Bessel potentials are expressed in terms of ${\bf F}_{\alpha}$ and $U_{\alpha}$, and on the deSitter and Anti-deSitter space we need closely related Gegenbauer functions instead.
@@
In our discussion we cover not only the massive and massless case, but also the tachyonic case. This case is quite curious, even though usually ignored in the physics literature. We also discuss identity (1.14), true for $m^{2}\geq 0$, but wrong in the tachyonic case. Remarkably, this identity sometimes, but not always, generalizes to curved spacetimes, as analyzed recently in [13].
In our treatment, we pay special attention to the distributional character of Bessel potentials. This is unproblematic in the Euclidean signature, where Bessel potentials are given by (locally) integrable functions. This is not the case in non-Euclidean signatures. In particular, it is interesting to look at the functions ${\bf F}_{\alpha}$ and $U_{\alpha}$ as defining distributions on the real line. With this interpretation in mind, well-known identities have to be reformulated, see e.g. (2.47).
Finally, let us mention that there exist a large literature about Green functions of the Klein-Gordon equation on curved spacetimes. In the generic context their explicit expression is not possible, and often instead of exact Green functions one restricts oneself to parametrices, that is inverses modulo smoothing terms. The existence of exactly four parametrices that generalize $G^{\mathrm{F}/\overline{\mathrm{F}}}$ and $G^{\vee/\wedge}$ is the result of a famous paper by Duistermaat and Hörmander [14]. It is also remarkable that expansions similar to (5.34)-(5.37) describe singular parts of these parametrices also in curved spacetimes, where they can be derived from the Hadamard recursion relations (see Chapter 4 of [15] or Chapter 2 of [16].) The universality of these singular parts is an important idea in Quantum Field Theory on curved spacetimes [7].
@@ 2.1 The ${}_{0}F_{1}$ equation
Our presentation of Bessel potentials will use extensively ${}_{0}F_{1}$ hypergeometric functions, closely related to functions from the Bessel family. Surprisingly, they are seldom used and discussed in the literature. Therefore, we devote this section to a concise exposition of their properties, mostly following [17] and [18]. In particular, we will treat these functions as distributions on the real line, as explained in section 2.5, which leads to useful distributional identities which we have not seen in the literature.
@@
$(z{\rm\partial}_{z}^{2}+c{\rm\partial}_{z}-1)v(z)=0.$ (2.1)
If $c\neq 0,-1,-2,\dots$, then the only solution of the ${}_{0}F_{1}$ equation equal to $1$ at $z=0$ is called the ${}_{0}F_{1}$ hypergeometric function:
$F(c;z):=\sum_{j=0}^{\infty}\frac{1}{(c)_{j}}\frac{z^{j}}{j!},$
where $(c)_{j}$ denotes the Pochhammer symbol:
$(a)_{0}=1,$
$(a)_{n}:=a(a+1)\dots(a+n-1),$ $n=1,2,\dots$
$(a)_{n}:=\frac{1}{(a-n)\dots(a-1)},$ $n=\dots,-2,-1.$
$F(c;z)$ is defined for $c\neq 0,-1,-2,\dots$. Sometimes it is more convenient to consider the function
@@
As $|z|\to\infty$ and $|\arg z|<2\pi-\epsilon$, $\epsilon>0$, we have
@@
$U_{\alpha}$ is the unique solution of (1.16) with this property. (Note that the validity of (2.7) extends beyond $|\arg z|<\pi$, that is, beyond the principal sheet of the Riemann surface.)
@@ 2.3 Relationship to Meijer G-functions
Solutions of hypergeometric equations ${}_{p}F_{q}$ can be expressed in terms of Meijer $G$-functions [19]. In particular, the ${}_{0}F_{1}$ equation can be solved by two distinguished functions
@@
Here, the contour $L_{1}$ goes from +$\infty$+ ∞ to +$\infty$+ ∞ and encircles ${\mathbb{N}}_{0}$, and the contour $L_{2}$ also goes from +$\infty$+ ∞ to +$\infty$+ ∞ and encircles ${\mathbb{N}}_{0}\cup({\mathbb{N}}_{0}-\alpha)$, both counterclockwise. Computing the residues and using the connection formula (2.8) we obtain
@@ 2.4 Relationship to Bessel functions
In the literature, the ${}_{0}F_{1}$ equation is seldom used. Much more frequent is the modified Bessel equation, which is equivalent to the ${}_{0}F_{1}$ equation. It is given by the operator
@@
Even more frequent is the (standard) Bessel equation given by:
@@
The function ${\bf F}_{\alpha}$ is also seldom used. Instead, one uses the modified Bessel function and, even more frequently, the Bessel function:
@@ 2.5 ${\bf F}_{\alpha}$ and $U_{\alpha}$ functions as distributions
where the right-hand side should be understood as the limit in the distributional sense. Note that for ${w}\neq 0$ these distributions are regular (in the sense of Appendix A.2) and given by analytic functions:
@@
(Note that both sides of (2.47) are zero for $w>0$). It is easy to see that for $\alpha\not\in{\mathbb{N}}$ (2.47) is a correct distributional identity, where the lhs is the product of the distribution $w_{-}^{-\alpha}$ and of the smooth function ${\bf F}_{-\alpha}(w)$, whereas the rhs is a linear combination of distributions defined in (2.42). (2.47) can be decomposed into a singular and regular part as follows: